OpenStudy (anonymous):

-((x-2)x/4)+3/4 = 0.75 solve for x

OpenStudy (anonymous):

-((x-2)x/4)+3/4 = 0.75 −0.25x2+0.5x+0.75=0.75 −0.25x2+0.5x=0 x=0,2

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np